40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 9 of 60  Not logged in ELOG logo
ID Date Author Type Category Subject
  2594   Mon Nov 30 11:28:30 2020 aaronNoise HuntingLaserpll

Entered lab about Mon Nov 30 10:21:35 2020, after taking a COVID test through Caltech's new surveillance testing program.

I'll pick up where Shruti left off on the beat note. The comb of sidebands becomes a single line  remains a comb when the PID is offKoji suggests maybe the (PLL) PID is oscillating at 10Mhz.

  • Turned on both laser drivers, TEC, 15V PD power supply, HP 8560E spectrum analyzer.
    • E laser TEC was set to 8.301 kOhm
    • W laser TEC was set to 8.992 kOhm
  • Located the thermistor resistance on the RIO data sheets
    • Diode #104978 (E path) has thermistor resistance 10.050 kOhm at the nominal setpoint (25 C)
    • Diode #104987 (W path) has thermistor resistance 10.940 kOhm at the nominal setpoint (23 C)
    • Oddly, Dmass' elog on unboxing these lasers lists a different operating temperature than the datasheets
  • Adjusted the PID on both TEC
    • First, turned down I and D to 0. Next, increased P until there was oscillation. Added back a little D and turned down P until no oscillation. Increase I until it moves to the setpoint.
    • Checked 'step response' when the laser is turned on, and adjusted PID until no overshoot or oscillating on the way to setpoint, and equilibrium reached within seconds. This did not affect the sidebands.
    • Adjusting P while moni
    • toring the beat spectrum results in no change in the sidebands, until the temperature starts to oscillate and the beat is lost.
  • Checking out the HF current mon on a 200MHz oscilloscope to see if something's going on there. The HF drive input is open, so there should be no peaks.
    • W laser has a line at 9.75 MHz, 20 dB above the noise floor.
    • E laser has a line at 22.5 MHz, 20 dB above the noise floor (on second look, this peak is not present. Lots of adjustment between the two measurements though, and I'm not sure what did the trick)
  • I'm seeing the forest of sidebands hopping side-to-side (~38 MHz hops). Modehopping? [no, it was a triggering issue] I adjust the TEC setpoint to the nominal values on the datasheet.
    • The hops are always by exactly 4 times the sideband spacing (~40MHz, so the second to the right sideband frequency at +20 MHz coincides with the second to the left sideband frequency at -20 MHz after the hop)
    • Hm... turning off both lasers, I see several noise peaks in the dark spectrum that 'turn on and off' every other refresh of the spectrum video...
      • This is present even when I take single spectra manually (not continuous). Every other spectrum has some noise forest around 20MHz. Is this an artifact of the SA settings? No reason something at MHz should exhibit such repeatable on/off behavior for my random-near-Hz button presses.
        Indeed, when I turn on the 'frequency counter' on the spectrum analyzer, the hopping goes away... why?
      • With frequency counter on, I now see two sets of sideband peaks. What's the spectrum analyzer doing for demodulation, this looks like some artifact. Perhaps some saturated stage of the spectrum analyzer is causing sidebands. Here's a nice technical note from HP on spectrum analyzers.
  • Larger changes of temperature and current moves the forest of peaks out to higher frequency (after being lost for intermediate values of T and I).
  • The forest of modes is still there even with both PID off.

 

exit Mon Nov 30 16:12:04 2020

Attachment 1: IMG_0195.jpg
IMG_0195.jpg
  2592   Mon Nov 30 10:26:17 2020 ranaPhotosSi fabSi cantilever photos

Thanks, the photos are now on the shared drive.

  2590   Sat Nov 28 21:53:12 2020 ranaPhotosSi fabSi cantilever photos

For storing lab photos in W Bridge, you can use our shared google acct instead so that we all have access to it (see chat for secrets)

  2589   Fri Nov 27 19:43:33 2020 KojiUpdateLab Workpsoma locking

Isn't the PID oscilating at 10MHz?

  2588   Fri Nov 27 13:38:18 2020 ShrutiUpdateLab Workpsoma locking

Attachment 1: Video of spectrum analyzer with zoomed out beat after turning off the PID loop of west laser

Attachment 2: Another image of the zoomed in spectrum when the PID is on.

Attachment 1: Beat.mp4
Attachment 2: BeatZoomed.pdf
BeatZoomed.pdf
  2587   Tue Nov 24 15:58:00 2020 aaronNotesEquipment Loanmoku CTN to Cryo

Anchal dropped off the Moku from CTN, along with its USBA->USBC cable, power cable, and ipad.

  2586   Tue Nov 24 13:23:09 2020 aaronPhotosSi fabSi cantilever photos

Entered lab around Tue Nov 24 13:24:57 2020 to finish photographing Zach's cantilevers.

some things about cameras, and in particular the FinePix F300 EXR

  • ISO -- the camera's sensitivity to light. More sensitive means more noise, but also more signal (useful when exposure time must remain short). 
  • aperture -- size of the opening before the lens. A wide aperture yields a shallow depth of field and lets in more light, but can cause blurriness in the foreground and background. Narrow aperture lets in less light and widens the depth of field, but can lead to diffraction effects or not enough exposure time depending on the application.
  • shutter speed -- how long with the shutter remain open? All the usual tradeoffs of integrating.
 
For these shots, I have the camera mounted on a tripod and close enough to the cantilevers that the subject takes up the full field. I've turned off the overhead lights and oriented the bright, fluorescent desk lamp away from the camera and slightly up. I'm reflecting some diffuse light back to the cantilevers with a large kim wipe (and my white face / lab coat). I've set minimum ISO, long exposure time, and large aperture. Since I'm handling nominally clean Si, I'm wearing gloves, lab coat, mask, and hair net. I covered the wall behind the shot with paper towel to provide a dark, uniform surface that isn't visible in reflections off the wafer containers and doesn't backlight the shot. The photos are not taken from above because doing so resulted in reflection from the top cover, which I wanted to keep on to avoid dust.
 
Looks like 5-7 cantilevers could be high Q (no visible contaminants, pitting, cracks, etc). I think there's also a pair of high Q cantilevers in Zach's cantilever cavities cryostat, and maybe one in the cantilever op lev cryostat (QIL). If we measure the Q of the most promising bare cantilevers, we can identify the 3ish best candidates for aSi coating.
 
Photos can be found in this album. I've pulled a representative good-looking cantilever and attach them here, along with the photo booth setup.
 
attachment 1: the photo booth
attachment 2: the lines near the top are reflections from the edge of the container. The long sides are parallel, and can give you a sense of the angle of the photograph. 
 
While doing this, I entered EE to retrieve a spare battery for the camera, and later again to return the camera to its place by the sink.
 
left around Tue Nov 24 17:50:48 2020
Attachment 1: DSCF3580.JPG
DSCF3580.JPG
Attachment 2: DSCF3572.JPG
DSCF3572.JPG
  2585   Tue Nov 24 10:16:04 2020 shrutiUpdateLab Workpsoma locking

Updates

- Found a suitable power cable M-M for the New Focus 0901 power supply on the east table (I did not realize yesterday that these were the same cables). Then I checked the voltage on the pins and they were fine.

- Using the New Focus 1611 (1 GHz PD) powered by the New Focus 0901 +-15 V / 0.3 A max. power supply, I tried finding the beat note. I looked at the RF output on a HP 8560 E spectrum analyzer and the DC output on an oscilloscope.

The DC output ranged from 500 mV to over 1 V as I scanned the temperature of one or both lasers.

- When the east laser temperature read roughly 8.34 kOhm and west was 9.04 kOhm I saw a pattern as in Attachment 2.

Changing the temperature slightly did cause the peaks to shift about, and further when I changed the polarization of the east laser using the HWP the height of the peaks varied. They also disappeared when either of the beams were blocked.

The estimated peak power in the taller peaks is ~0.1 µW from the plot.

- I also tried scanning the temperature of both lasers again to possibly find a single peak. No luck yet.

Today, I didn't check the alignment very carefully and I probably have to tune it further after the changes that Aaron and I made over the past few days.

The next step is to do the phase-locking.

 

 

Attachment 1: BeatSetup.pdf
BeatSetup.pdf
Attachment 2: BeatOrNot.pdf
BeatOrNot.pdf
  2584   Fri Nov 20 19:24:00 2020 aaronElectronicsElectronicspower cables

Entered lab Fri Nov 20 19:24:32 2020, usual sanitation.

[blue "Photodiode Power Supply] Looking for a DB9 to BNC adapter. I found this spider instead -- close enough. Use multimeter to measure 24V between pints 4 an 9... not promising. Confirm power is connected, no signal on the frontpanel BNCs either. Could remove this one and take a look on the benchtop, but above is...

[Newfocus +-15V current-limited power supply] Has 3 pole bananas and a power switch on the front. Found a power cable for the back. There's a bananas to 3-pin LEMO already there. Double check the voltage with a multimeter. Alas the connector doesn't fit the PD, but should be some cables in the EE shop or elsewhere...

Didn't find the right connector in EE. On the 'power cables' rack (NE corner Cryo), there was a M-F connector, but I need M-M. Could cannibalize the 12VDC supply? I think for now +-12V is working, so should look a bit more.

 

I moved the power cables for our preamps for better strain relief (attachment 2 is the before photo). 

I also had left this ND filter sitting on the table (attachment 1). Yikes!

More photos here.

ExitFri Nov 20 20:39:43 2020

Attachment 1: DSCF3541.JPG
DSCF3541.JPG
Attachment 2: DSCF3542.JPG
DSCF3542.JPG
  2583   Thu Nov 19 16:34:01 2020 aaronDailyProgressGeneralcantilever photos, laser intensity noise

Entered lab Thu Nov 19 16:33:57 2020. Usual sanitation, personal reminder to report campus access with Caltech.

Cantilevers

we want to coat some of Zach's cantilevers with a-Si so we can make a cold Q measurement. I've started to take some photos, but have become tired and will finish tomorrow. There are O(5) suitable cantilevers produced in January 2018, but I'll have to dig a bit more (or ask Zach) to determine what's what. We can measure the Q of the most promising few cantilevers to be sure they're acceptable.

I borrowed the digital camera from EE shop, but left its case (which is very dusty).

Laser intensity noise

Sending W path beam to a Newfocus 1811 to measure free running laser intensity noise.

Following Shruti's recent diagram, I moved the Newfocus 1811 into position after OMTL1. I also moved PO1.1 back into the beam path, so I can use it to align into the 1811. Turn on the E laser and TEC, also had to move Ma for alignment. I still don't have a +-15V power supply, will ask around. Turn off the laser and TEC before exit at Thu Nov 19 20:41:22 2020

  2582   Thu Nov 19 11:23:44 2020 shrutiUpdateLab Workpsoma locking

Attachment 1: An updated version of the diagram in elog 2577 where the path lengths to the beat beam-splitter are identical. The fiber launchers and some components have been moved around, but everything after PO1.1 along the beam has been retained as before.

Attachment 2: Retaining the same configuration to the beat BS, the cavity with Mach-Zehnder interferometer has been added. Also the path lengths to the MZ input BS along both laser beam paths have the same length. Except for the ring cavity, the Mach-Zehnder is also balanced.

Attachment 3: Updates pertaining to the current setup

  • Work in progress to achieve the configuration in Attachment 1.
  • I have switched the two PDs so the beat can be measured with the Newfocus 1611 (has a larger BW) and the noise measurement with the east laser can be done with the Newfocus 1811.
  • I removed the 10/90 pick-off circled in green, so that would have to be added to the path to continue the noise measurement.

 

Attachment 1: PLL_FS_sym.pdf
PLL_FS_sym.pdf
Attachment 2: MZ_PLL_FS_sym.pdf
MZ_PLL_FS_sym.pdf
Attachment 3: NewBeat.pdf
NewBeat.pdf
  2581   Wed Nov 18 08:24:02 2020 shrutiUpdatePSOMAtemp control and transverse beam profiles

D [in.]

West laser (X) [um]  West laser (Y) [um] East laser (X) [um] East laser (Y) [um]
2 327.2, 336.5 332.0, 340.9 327.1, 332.8 330.0, 335.2
3 358.2, 269.6 363.4, 374.7    
4 420.4, 439.3 417.8, 439.8    
5 511.7, 549.2 510.2, 549.9 527.6, 550.7 519.6, 540.4
6 642.8, 688.0 630.6, 674.9    
7 766.4, 807.9 754.6, 801.4 778.0, 848.3 716.0, 782.7
8 891.1, 932.6 894.7, 942.1    
9 974.3, 1023.1 943.9, 1000.6    
10 1142.6, 1193.9 1152.7, 1203.5 1092.2, 1159.7 1040.6, 1103.4

 

This is the data using the Data Ray Beam'R2 profiler with the InGaAs window. Attachment 1 contains images of each of those profiles.

D: distance from fiber launcher in inches; The two values in each of the cells are [Clip 13.5%, 4 sigma] respectively, i.e., the method used to calculate the beam widths.

The previous measurement using a razor blade refers to 'sigma' which I believe explains why these values are 4 times larger.

These profiles were taken with temperature stabilized such that the powers were ~1 mW.

East laser set to 8.070 k Ohm, West laser set to 9.065 k Ohm. I don't understand why there is such a difference.

 

Other updates:

I had hooked up the ITC 502 combi controller to the west Rio laser and used only its temperature controller. (I believe both the thermistors that measure the diode temperatures are TH-20k Ohm.)

Both the PID controllers work satisfactorily: the TED 200 C with the east laser stabilizes to within few Ohms of the setpoint thermistor resistance within some seconds;

the ITC 502 stabilizes at a similar rate but at an offset of ~10 Ohms despite the integrator being set to maximum. I fiddled around with the P and I settings a little but realized that this configuration seemed optimal.

 

To measure these profiles at different distances I moved the fiber launcher head and then replaced it back to its original position, roughly.

Attachment 1: beamprofdata.zip
  2580   Tue Nov 17 14:56:06 2020 aaronDailyProgressGenerallab entry

I entered the lab somewhat before Tue Nov 17 14:56:22 2020. Exited Tue Nov 17 16:39:12 2020

Hand sanitizer on entry, also sanitized the bulky green laser goggles before and after my use (forgot contacts). Turned off the laser, sat at desk and considered turning on the laser. Took a break. On my walk I wrote this haiku

Took another break.
Why align these mirrors when
photons have no mass?
 
I'm going to this seminar, so that's all for today. If nothing else, a reminder (even to myself) to always elog, no matter how small. 
  2579   Mon Nov 9 14:40:46 2020 shrutiMiscEquipment LoanStuff for N2 transfer

I've placed the following items outside the cryo lab:

1. Cryo liquid N2 dewar

2. Funnel

3. Tube

4. Two pairs of cryo gloves

Attachment 1: IMG_0334.pdf
IMG_0334.pdf
  2578   Wed Oct 28 18:10:27 2020 aaronNoise HuntingLaser

Lab entry

in: Wed Oct 28 18:11:00 2020
out: Wed Oct 28 21:05:58 2020
 

E laser noise

  • Moved the Newfocus 1611 from W path cryo cantilevers transmission, to just after the PSOMA beam launch.
    • nominal AC gain: 700 V/A
    • nominal DC gain: 10 kV/A
    • output impedance: 50 Ohm
    • Max input power: 1 mW DC (10 mW CW)
  • Measuring the following spectra across the PD band (~30kHz-1GHz) by connecting the 1611 AC output directly to an HP 8560E (50 Ohm input). I'm monitoring DC with TDS3024B oscilloscope (1 MOhm input).
    • dark current
    • free running E path laser intensity noise

Optical layout: beam launch -> lambda/2 -> steering mirrors -> lens 1 -> ND 0.6 -> lens 2 -> PD 1611

There is only a 12 VDC power supply compatible with the 1611 power port, but the PD requires +- 15V. Surely there's one somewhere. Perhaps this is why I observe only -6 V on the DC mon with 1mW input power at 1550nm (checked against the Thorlabs S122C; I expected -10V). Maybe the beam is too large.

  2577   Wed Oct 28 12:42:31 2020 shrutiDailyProgressLab Workpsoma locking

Phase-locking the two lasers:

Updates:

- Although when we talked about adjusting the MZ-phase, we decided that having the phase/path length control with fiber components might be better initially (Refer Attachment 2), for now I began doing everything in free-space.

- Attachment 1 shows the setup as it is now. Previously I'd placed polarizing beam splitters instead of 90/10 beam-splitters because I thought it would be easier to work with, but now changed my mind and decided to stick with what we planned.

Next steps:

(Once the beat is obtained on the spectrum analyzer)

1. Adjust set-point temperatures to adjust beat frequency to the right frequency.
Since ideally we want the two frequencies to be identical, it might be a good idea to add an AOM to one of the paths so that the first order beam is mixed with the other laser and this beat be compared to a stable reference for phase-locking.
But, initially we plan to skip the AOM.
2. Adjust the PID parameters if needed
3. Add electronic components

- Measure the laser frequency noise


29 Oct 20:

I've added Attachment 3 -- which is the current free space version and some PLL electronics. 

- It does not show the Mach-Zehnder part as that will be added only later

- This setup is asymmetric but in a future version we will change that

Attachment 1: Setup2020Oct27.pdf
Setup2020Oct27.pdf
Attachment 2: psoma_PLL.pdf
psoma_PLL.pdf
Attachment 3: psoma_pll_freespace_intermediate.pdf
psoma_pll_freespace_intermediate.pdf
  2576   Tue Oct 27 15:57:27 2020 shrutiDailyProgressLab Workpsoma locking

Today I modified the optical setup with the aim of obtaining the beat between the two diode lasers for phase-locking.

I added pick-off polarizing beamsplitters with HWPs in each path for now (to be able to adjust their power) and mixed them at a 50/50 non-polarizing beam-splitter to eventually reach a Newfocus 1811 low noise PD.

I will add pictures and more details later.

 

  2575   Fri Oct 23 19:06:05 2020 gautamUpdateEquipment LoanTwo resonant RFPDs ---> 40m

I was in the Cryo lab between 1215-1230 this afternoon. I removed two resonant RFPDs from what was Johannes' setup (encircled in Attachment #1). I also brought a SR554 preamplifier to the 40m.

I was wearing the usual PPE (gloves, face mask) while I was in the lab.

Attachment 1: IMG_8903.JPG
IMG_8903.JPG
  2574   Thu Oct 15 22:15:11 2020 Ian MacMillanComputingSimulationModeringer Simulation

This is the spectrum coming off of the sample. there should be a peak at 1038Hz... but there isn't. And what is even weirder is that the spectrum analyzer that is built into Simulink shows a peak where I expect but when I do it here it doesn't show up. 

Update: I think I have found why there is a discrepancy between the two versions of the power spectrum. The spectrum analyzer in the Simulink model requires non-continuous data so you have to use a block to make your data discrete. The sampling rate of that block affects where the peak of the mode is seen. So it seems that that the mode seen in the previous post was just caused by making the data non-continuous.

 

Attachment 1: ModeRingerSpec.pdf
ModeRingerSpec.pdf
  2573   Thu Oct 8 00:31:11 2020 Ian MacMillanComputingSimulationModeringer Simulation

I got the script to run the simulation and added color to the diagram just to make it pretty :0

The next step is quantifying the error of the loop. My plan for this was to just calculate the Q from figure 3 of the moderinger paper. Then I can see if that value is consistent. Aaron suggested settling time and phase margin of the loop. So that is the next step (once I figure out how to do that)

Also I added the code to the Qryo github

Attachment 1: ModeRingerLoopDig.pdf
ModeRingerLoopDig.pdf
Attachment 2: ModeRingerSim.zip
  2572   Tue Sep 29 11:00:53 2020 aaronComputingSimulationModeringer Simulation

Beautiful! Want to push this to the repo under git large file storage?

Quote:

The step function represents the excitation of the sample. Ideally, it would excite to the setpoint and stay there but for some reason, it is jumping way past the point before returning. By messing with the gain and the frequency of the low pass filter I could get a variety of results the best is shown below at 50 gain and 40 rad/s.

The overshoot is interesting! To understand the loop shaping, I suggest checking out Gardner's Phaselock Techniques or Astrom and Murray's Feedback SystemsThey both have sections on optimal PID controller design (at least Astrom and Murray do). You can make a pole-zero plot to help choose the location of poles and zeros in your loop shaping filter (the discrete zero-pole TF we added, after the gain).

 

  2571   Tue Sep 29 02:02:54 2020 Ian MacMillanComputingSimulationModeringer Simulation

Working with Aaron's suggestions (In the previous post) we got the mode ringer to converge. Previously the loop would continue to excite the sample to infinity but by fixing the following things we were able to get the step function to converge.

  • Added a discrete zero-pole transfer function after the 100 gain
  • Included the RMS block
  • Disabled zero-crossing detection in the saturation
  • Fixed the sign of the s^2 in the bottom of the plant TF (the i^2 included in s cancels the negative)

The step function represents the excitation of the sample. Ideally, it would excite to the setpoint and stay there but for some reason, it is jumping way past the point before returning. By messing with the gain and the frequency of the low pass filter I could get a variety of results the best is shown below at 50 gain and 40 rad/s.

The power spectrum shown is taken from the spectrum analyzer shown in the loop. It shows what I would expect with a peak around our mode frequency of 1038 Hz.

Next Steps:

  • Figure out how to run the model entirely from a script. (I got this partially working but it is not fully functional)
  • Fix the step function it still seems to be wild and we NEED to fix that before making meaningful progress on this project.
  • Add some sort of error estimation: quantify how good this loop is at measuring what we want.
    • How can we optimize our loop parameters to minimize this error?
  • Make it look pretty (black and white is gross)
Attachment 1: ModeRingerLoop.pdf
ModeRingerLoop.pdf
Attachment 2: StepPlot.pdf
StepPlot.pdf
Attachment 3: PowerSpectrum.pdf
PowerSpectrum.pdf
Attachment 4: ModeRingerSim.zip
  2570   Tue Sep 22 21:30:07 2020 aaronDailyProgressLab Worklab entry, supplies

notice of lab entry

Tue Sep 22 21:33:30 2020

inventory

I'm cleaning a bit, and gathering items not in use or in need of repair. They would make less mess in my office.

took an inventory of optics cleaning supplies, first aid kit, general cleaning supplies, wipes, etc. I found most were included in the first round inventory, but I took photos this time to convince my future self of object permanence. Will add to the wiki and update in a bit.

entered QIL, CTN

gloved up, shoe covers, went to QIL to check out the sprinklers and CTN to grab a GHz spectrum analyzer (HP8560E).

out: Wed Sep 23 00:20:03 2020

  2569   Tue Sep 15 12:29:41 2020 ranaLab InfrastructureHVACHEPA filters for PSOMA enclosure

I suggest just using what's used for th QIL table.

  2568   Thu Sep 10 10:21:34 2020 aaronLab InfrastructureHVACHEPA filters for PSOMA enclosure

Yesterday into today, I've been shopping for laminar flow HEPA fan filter modules for the PSOMA optical enclosure. I didn't find a lot of LIGO documentation listing specific filters, but here's what I've found online with some downselection on 'low vibration / high filter quality'. Please let me know If there's a company we often use, or if you can help direct this search at all. 

Another consideration is flow rate relative to our volume. I can do this calculation, but what is the particle density (eg at 0.3 um) we want to achieve inside the enclosure? I realize that I never got the particle counter recording despite it being on my list, so I'll try to do that remotely today. We need to know the ambient particle count and the clean volume in the enclosure.

Company

Part number / name

Price

Filter

Sound

Link for info

Notes

Terra Universal Smart Whisperflow Fan Filter Unit $1,111 HEPA (H10-100k) 40 dB @ 90 ft/min velocity @ face datasheet, product page The 'Smart' model has 9dB less noise than the standard, but costs about $500 more. Also comes in 'low profile' version.
Clean Air Products CAP118   HEPA -- product page requested quote
Thomas Scientific Fan filter module, "low noise" $1,760 HEPA   product page  
Price Industries LFDC series laminar flow diffusers       product page, datasheet  
Envirco Mac 10 original     51 dBA @ 90 FPM datasheet referenced in T2000382
A-J MFG SSLFHFD-FP "criti-clean" fan filter   HEPA 45dBA @ 90 FPM product page claims to be the lowest-height unit on the market (13.7''). The RFQ page looked fishy though, so I held off.
FungiFun (website)       link lists a number of other filter suppliers, by country

*all prices listed here are from publicly available pages

  2567   Wed Sep 2 16:46:39 2020 aaronThings to BuyDrawingsdimension-full optical layout

I updated the stage 1/2 optical layout to be more detailed after getting a sense of the sizes of things again last week. Even though this isn't how the table is currently set up, it might be good to accommodate future vacuum chambers in our earlier designs to minimize how much we need to move and realign optics. 

  • just for convenience, I gave all of the vacuum chambers the footprint of the gyroscope vacuum chambers. I chose this footprint rather than the cantilever cavity cryostat footprint to allow 45 deg aoi on the mach zehnder BSs (the cryostat windows are all parallel)
  • The chamber containing the output BS and BHD is cramped. If these all need to be in the same vacuum, we should get a larger vacuum; if the BHD and LO path can be moved to air that works too.
  • If output BS chamber is larger and input MMT allows, the entire MZ can be shifted to the right (north). 
  • Picked off LO from input BS AR surface
  • Need additional mode matching lenses into some of the photodiodes -- certainly into PL_RFPD, but I think also into the other PDs?
  • The beam to TRANS_DCPD is probably geometrically unphysical. 
  • The cavity length is limited by the size of the cryostat.
  • Perhaps want free-space EOM, EOAM, Faraday isolator?
  • Not sure I've dumped all beams properly, for example the reflection from BHD_PD1 and from BHD_BS AR surface.

I will update the PSOMA hardware inventory tomorrow to reflect the additional details in the new drawing. The updated diagram is available on git LFS, and the hardware inventory now reflects the diagram up through stage 1.

Attachment 1: psoma_stage2.pdf
psoma_stage2.pdf
  2566   Wed Sep 2 16:19:57 2020 shrutiUpdateLab Worklab entry, information gathering

Notice of lab entry: 20 Sep 2020   evening

Fiber modulators on the table :

1. Intensity modulators (BW: up to 12 GHz) MXAN-LN-10

2. EOM phase modulators (BW: up to 150 MHz) MPX-LN-01

Dimensions of vacuum cans mentioned in attachments.

Attachment 1: 7BA40146-4F86-4819-8D6C-FFDC8F246E50.jpeg
7BA40146-4F86-4819-8D6C-FFDC8F246E50.jpeg
Attachment 2: 129FE089-1E57-4E6D-9E77-F80F8E124CD1.jpeg
129FE089-1E57-4E6D-9E77-F80F8E124CD1.jpeg
  2565   Mon Aug 31 10:00:25 2020 aaronUpdateLab Workoptics setup

Here's the layout.

Some easy things that should be changed:

  • Transmission through the cavity end mirror should go to a steering mirror before PDA20CS
  • Should clamp the cables (eg power to PDA255) with something soft so they stay in place and aren't strained
  • Need to dump the reflection from transmon PD. Do we really need to dump transmission through mirrors with frosted backs?
Attachment 1: 200828_layout.jpg
200828_layout.jpg
  2564   Fri Aug 28 16:59:59 2020 aaronUpdateLab Workoptics setup

I measured the transmission of the Coastline 1m mirror at 180. ppm (S122C).

Alignment procedure while setting location of optics:

  1. use 2x irises to constrain a beam path at the locations of the eventual MZ input BS and the cavity BS.
  2. Place the first (R~1) flat mirror at 45 degrees centered on the beam line. Use this mirror to steer into an iris at the location of the cavity's curved mirror.
  3. Use steering optics to direct the eventual transmission beam into the trans mon PD (PDA 20CS)
  4. Place the curved mirror at near normal incidence. Use this mirror to steer the beam through irises at the location of the MZ output BS and cavity BS.
  5. Use steering mirrors to send the beam to the refl PD (PDA255).
  6. Place the cavity BS and steer the reflected beam to the refl PD.

Alignment procedure subsequently:

  1. Misalign the cavity BS
  2. Use the cavity mirrors to steer the circulated reflection into refl PD.
  3. Use cavity BS to align prompt reflection into refl PD.

 

  2563   Thu Aug 27 13:24:30 2020 aaronLab Infrastructurestuff happenslocked out of optomechanics cabinet

Not sure how, but none of the drawers of the blue optomechanics cabinet are opening. I don't have a key. Here's what happened

  1. I closed one drawer ("A") and opened the drawer below it ("B")
  2. Closed B
  3. Opened A and B simultaneously; I could not open only B, only A and B together. Close A and B.
  4. Open both A and B. Close only A.
  5. Close B
  6. None of the drawers now open, even A and B together. seems locked

Found someone who's had this problem before, might give it a try...

[update]

This worked, I used the metal meter stick to unlock the drawer.

[Stephen's updates]

--> note that link formatting breaks link for me, so here it is - https://www.practicalmachinist.com/vb/general/help-my-lista-locked-me-out-how-do-i-open-201606/

--> wrote up a similar experience with additional detail ENG_Labs/260

  2562   Wed Aug 26 12:12:22 2020 aaronUpdateLab Workoptics setup
  • aligned beam along the NS axis using two irises and the existing 2x f=100 lenses from the E path.
    • HWP->steering mirror-> 90-10 BS -> iris -> lens -> lens -> iris -> PD
    • all transmitted beams dumped
  • Mounted a 1'' x 1m FL Si mirror in a polaris mount. Made a ring cavity with that and 2x mirrors coated for 45 deg.

optics

Can anyone tell me the specs / history of some of the custom optics in cryo? I'm mounting the 1m Coastline mirror and will start with that in the PSOMA cavity.

  • Laseroptik 1'' mirrors. Batch # 19028Kel. Part(?) numbers L-13997, L-13998. (these are in red boxes cases w clear lid)
  • Coastline Optics 1.0''x0.25'' FS substrates. Coated for 1550nm at 45 degrees. (these are in stacks cylindrical clear cases, and include some witness and rejected samples)
  • Coastline optics 1'' silicon substrate, 1 meter focal length. SN 1.0-Si-1.0M
  • Photon Laseroptik 1'' FS substrates (~50), labeled 75S308926. Also labeled with a wedge angle or what looks like a focal length (no unit)
  • Two stacks from Gooch & Housego, taped in bubble wrap with the data sheet. These are coated for 1550nm, transmission ~1% S and ~5% P at 45 deg.
  2561   Tue Aug 25 13:42:40 2020 aaronUpdateLab Workoptics setup

setting up PSOMA beam path

I took some photos of the existing layout. I'll just take apart the E beam path, and leave the W path unchanged for now as reference.

I moved the E fiber output coupler closer to the edge of the table, to make this path easier to reach.

Hopped around on the laser hysterisis curve for a minute. To optimize the temperature,

  1. Started with TEC on near room temperature, then turned on the laser driver with the current set to maximize power on the meter (S122C)
  2. increased the temperature setpoint (decrease resistance setpoint) slowly until the laser power starts to decrease.
  3. Turn off the laser driver. Decrease temperature setpoint to just below the maximum power setpoint.
  4. Turn on the laser driver. Total power is now 4.6mW, compared to ~2mW with TEC off.
  2560   Tue Aug 25 13:23:49 2020 aaronThings to BuyGeneralcryo lab inventory

I started a cryo lab inventory that is separate from the PSOMA hardware inventory, and intended for stock items in the lab (optics, electronics, clamps, general safety and cleaning supplies, etc). It will be a work in progress. Both are accessible to anyone logged in to google drive with their ligo.org credentials.

Both are also linked on the PSOMA project wiki.

  2559   Mon Aug 24 13:22:32 2020 aaronUpdateLab WorkalaPSOMA

Did some mode matching, see the git.

  2558   Mon Aug 24 13:21:40 2020 aaronUpdateLab Worknotice of entry

Notice of lab entry

Date of entry: 24 Aug 2020

time of entry: 12pm
time of exit:  
 
Purpose: stage 1 cavity setup
  2557   Wed Aug 19 12:15:40 2020 aaronUpdateLab WorkNotice of entry

The cameras were unfortunately lost in the mail, but we can use my laptop or other camera. Ended up leaving to do a couple comsol things that needed completing today.

  2556   Wed Aug 19 12:11:47 2020 aaronUpdateLab WorkNotice of entry

Notice of lab entry

Date of entry: 18 Aug 2020

time of entry: 12pm
time of exit:  
 
Purpose: Identify some combination of optics that will let us mode match to our ring cavity, set up PDH locking with the existing ~33 MHz modulation.
  2555   Tue Aug 18 21:35:49 2020 aaronUpdateLab Worknotice of lab entry
I'm taking inventory of our optics supply, in between figuring out what's wrong with the temperature log channels.

The inventory is in Clickup, which is a new organizational tool I'm trying out. There' an easy csv export, so I can get it elsewhere if/when we want. We have a wide variety of lenses:

Focal length (mm) 25.4 38.1 100 103.22 154.82 200 225 250 258.04 309.65 412.86 500 850
Quantity @ 1550nm 1 4 2 7 2 2 6 2 1 3 12 1 2

I found fewer curved mirrors, but there were a couple.

Steps I took for the temperature sensor:

1. Tried to see what's the temperature by reading the current temperature in ndscope or dataviewer. In dataviewer, go to 'Signal' tab, and enter the channel name or find it on the list of slow channels. For ndscope...

nsdscope X1:AUX-LAB_TEMP_F 

No data appears on either. I restart cymac1, which seemed frozen, but still nothing.

2. What's going on with these channels?

$ caget X1:AUX-LAB_TEMP_F
X1:AUX-LAB_TEMP_F     -459.7
$ caget X1:AUX-LAB_TEMP_K
X1:AUX-LAB_TEMP_K     -0.0166677

Looks like the channel is reading zero.

3. I traced the cable from the particle counter and found that it sends data to cominaux, the common auxiliary machine for the lab.

ssh cominauxThat's a little surprising to me, because
cd ~/services/modbusIOC/
less CRYOXT.db

This is the database file that defines the channels on cominaux. I search for 'LAB_TEMP_F' and find the epics record for the temperature channels. The epics records are all "calc" records, and the temperature in Kelvin is taken from X1:AUX-ACROXT_AI_15. This corresponds to channel 15 of the acromag slow ADC.

$ caget X1:AUX-ACROXT_AI_15
X1:AUX-ACROXT_AI_15     -0.0003333

 That's starting to make sense, the cable from the particle counter didn't go to the acromag ADC. Starting from the ADC channel 15, I traced the cable back to what used to be the AD590 temperature transducer.

4. Where did the IC temperature sensor go? Searching the elog and my dusty memory... neither readily recalls where it went. Let's get another one, they are cheap and easy to use.

  2554   Tue Aug 18 15:05:33 2020 aaronUpdateLab Worknotice of lab entry

Notice of lab entry

Date of entry: 18 Aug 2020
time of entry: 2pm
time of exit:   11pm

Purpose: Inventory lens supply, identify some combination of optics that will let us mode match to our ring cavity. Picked up some books on silicon from the library.

  2553   Mon Aug 17 15:52:13 2020 aaronDailyProgressLab WorkNotice of lab entry

Measurements around table

  • Clearance of lights -- 100''
  • clearance of  beam -- 105''
  • distance between lights -- 6'
  • there's also a sprinkler head a few inches lower than the lights
  • distance from top of table to ground -- 35''
  • distance from bottom of table to ground -- 24''
  • distance between sprinklers -- 8'

Could probably move the table a few inches from the wall and make use of the space between the lights for the enclosure. There also isn't much room in the back corner in the NS direction, and we may want to shift in either direction. Orientation seems as good as it could be. The ceiling above PSOMA is lower than above cryo cavs.
Will mark up photos and post.

 

Attachment 1: Ceiling_lookingSouth.jpeg
Ceiling_lookingSouth.jpeg
Attachment 2: Ceiling_lookingNorth.jpeg
Ceiling_lookingNorth.jpeg
Attachment 3: Clearance.jpeg
Clearance.jpeg
  2552   Mon Aug 17 15:10:58 2020 aaronDailyProgressLab WorkNotice of lab entry

Notice of lab entry

Date of entry: 17 Aug 2020
time of entry: 3:30pm
time of exit:   
 
Purpose: Make sure unnecessary electronics in the lab are off, due to the CA energy Flex Alert. Take some photos of cryo cavities enclosure and filter to assist with new enclosure design.
  2551   Fri Aug 7 11:35:09 2020 aaronUpdateLab Worklab work

Entered the lab at 11:00 am. The lab is far too hot (78F) and humid (45%)! 

Inventoried available photodiodes in the cryo lab, on the PSOMA wiki.

Fri Aug 7 12:55:53 2020
Facilities came to check the fire alarm. I let them in with shoe covers, they held a magnet up to the fire alarm and left. Total operation took less than 5 min.
[lunch and a couple meetings 1 - 3]

Pretty tired honestly, I submitted an order for a few things we want in the lab, such as:

  • 2x "ultra quiet" fans to circulate air
  • longer ethernet cable for reaching far end of PSOMA table
  • headset web cam so we can take video while doing lab activities
  • more plastic storage bins
  • food and water for our emergency kit

The PSOMA optics table is from TMC vibration control (TMC 784-29701-01). I sent them an inquiry for the datasheet, and also asked whether they sell clean enclosures or have application notes that might help us put one together. Also talked with Arian Jadbabaie a bit about the Hutzler lab enclosures, and got some photos.

Exit ~430 pm

  2550   Wed Aug 5 19:03:56 2020 aaronUpdateLab WorkNotice of lab entry

Notice of lab entry

Date of entry: 7 Aug 2020
time of entry:  11:00 
time of exit:   
 
edit: postponed this work from 6 Aug to 7 Aug.

Purpose: Measure optical table dimensions, start to assess what we need for an enclosure. Set up EOM/EOAM to take some transfer functions. Set up realtime model. Figure out why particle counter isn't logging to frames. Turn on the HEPA for cryo cavs table.

  2549   Tue Aug 4 14:08:44 2020 shrutiDailyProgressLab WorkInitial experiments and prep: Beam profile

Notice of lab entry: I will be at the cryo lab for a few hours today (04 Aug 20) afternoon


Update:

Beam profile results [razor-blade method]

At a few locations along the beam I measured the vertical and horizontal beam radii by measuring the power at different positions of blade edge across the beam. The power measured at the photo-diode was fitted with  a + b \times {\rm erf}(\sqrt{2}(x-x_0)/w_z)) where w_z is the beam radius at the location along the beam and x was either the horizontal or vertical position of the blade.

Results:

The standard deviation as estimated from the fits are lower than \pm0.002 mm for all estimated radii.

But with the crude nature of the setup and not having the laser temperature stabilized (resulting in power drifts while taking readings), I guess that the error in each measurement is higher. The data in Attachment 3 also has recorded the resistance of the temperature sensor at different points while taking the data.

Beam radii
Distance [in.] Horizontal radius [mm] Vertical radius [mm]
5 \pm 0.1 0.120 0.121
6 \pm 0.1 0.146 0.150
7 \pm 0.1 0.175 0.173
8 \pm 0.1 0.203 0.200
9 \pm 0.1 0.231 0.229

 

- Attachment 1 has the updated setup with the clamped translation stage

- Attachment 2 shows the measured points with the error function fits. The offsets between the different curves along the x-axis are arbitrary.

- Attachment 3 shows a linear fit and an estimated divergence angle. I've assumed that all points I measured are outside the Rayleigh range, i.e., away the beam waist. The caption '5H' means that this set of data was taken 5 inches away from the output coupler (fibre-to-free space) moving the blade in the horizontal direction.

- Attachment 4 has all the data and jupyter notebook with analysis

- Attachment 5 shows the beam spot on the view card

Attachment 1: IMG_0268.jpg
IMG_0268.jpg
Attachment 2: Erf.pdf
Erf.pdf
Attachment 3: Div.pdf
Div.pdf
Attachment 4: beamprofdata.zip
Attachment 5: IMG_0258.jpg
IMG_0258.jpg
  2548   Mon Aug 3 13:10:46 2020 shrutiDailyProgressLab WorkInitial experiments and prep

Notice of lab entry: I will be at the cryo this (03 Aug 20) afternoon

Lights fixed, safety, cleanliness:

- On entering the lab I noticed that the work with the light fixtures was completed. Unfortunately I could not avert it or get it all covered in time; I had assumed that I would be contacted beforehand but was not. But, by inspecting the table I do not think it looked any dustier than before. For any such activity in the future after we’ve cleaned the optics, I will remember to get it covered beforehand.

- The particle counter did not seem to be saving any data so I’m unsure what the effect of this activity was. The 1 micron particle count:
When I entered (1 pm): 80
(6 pm) : 70
This change might probably be entirely attributed to the air scrubber.

- When Aaron and I chatted with Calum today about COVID safety, Calum pointed out that turning on any HEPA filters in the lab would reduce the time between single-person occupancy to almost 0 for our room, as it would serve as an additional air scrubber. I tried locating a switch on the cryo cavity table for the HEPA fans but could not locate it today.

Laser and beam profile related:

- After reading through bits of the laser operating manual and turning on the east cavity laser, I turned on the TEC (thermo-electric cooler) box [TED 200 C] with neither the TEC nor the servo on. I used this just to read off the resistance values of the sensor. With the current drive off it was 11.2 k-ohms [22 C].

- I waited until the sensor readings stabilized to changing at the level of a few ohms around 8.5 k-ohms [28 C] in a minute and began measuring the beam profile. (See Attachment 1 for setup)
I could not mount properly the only translation stage I found since the holes on it were smaller than 8-32, and moreover, I could not locate anything suitable that fits into the 1/4" holes on the table with this.

- Nonetheless, when I took some data and plotted it, it looked like an error function so I took more data. Will post plots later after fitting and analyzing. I may have to repeat this again.

- The sensor for the power measurement was S122C (Germanium) with a range of 700-1800 nm and 40 mW. After setting the power meter to 1550 nm, the measured power of the entire cross-section was ~2.3 mW.

- Also in the image in Attachment 1, the two optics (HWP and mirror) near the fiber-free space coupler, were previously in the path of the laser and now moved to either side of the beam. No other optics were moved from the previous set-up.

Attachment 1: IMG_633BBA57553C-1.jpeg
IMG_633BBA57553C-1.jpeg
  2547   Fri Jul 31 08:51:00 2020 shrutiDailyProgressLab Workbeam profile prep

Notice of lab entry: I will be at the cryo lab today (31 Jul 20) 9 am - 2 pm.

Purpose: Look around for components [power meter charger, power connectors for PDs, flashlights, etc], laser operating manual; find out specifications of available mirrors, RF PDs

Also, there seems to be only one laser temperature controller in the lab. The west cavity laser TEC port is not hooked up to anything. For the time being this is okay since we'd be using only one laser.

  2546   Thu Jul 30 09:13:14 2020 shrutiDailyProgressLab Workbeam profile prep

Notice of lab entry: I will be at the cryo lab today (30 Jul 20) 9am-5pm.

Purpose: Beam profile the Rio planex laser

- Turned on the laser that goes into the east cavity on. Turned on the Tenma supply, then the laser current driver  D1500207 labeled'E'. Saw a tiny, but bright, green spot on the detector card.

- Located razor blade and translation stage, for the beam profile measurement but not sure which power meter/ photodiode to use for 1550 nm. I plan to move the optics on the path of the east laser (to the east cavity) in order to do this measurement and later set up the initial ring cavity

- Aaron suggested I do another airflow measurement within the lab. Without the scrubber, there seems to be no measurable flow around either the Cryo-Q table or the cantilever table. I moved the scrubber into the lab near the workspace at the north side of the lab and let it run on medium speed. A few cm from the scrubber it the airflow rate is almost 5000 CFM, but it becomes <40 CFM even as close as the Cryo-Q table and again becomes negligible near the cantilever table.

- I set up the translation stage, moved relevant optics, but will continue to measurement later after I've researched and located a 1550 nm power meter

Two of the four overhead lights over the west (cantilever) table don't work anymore. Three other overhead lights are very dim/out.

 

 

  2545   Tue Jul 28 16:13:56 2020 aaronUpdateLab WorkAir flow measurement

Conclusions

  • I propose we continue scheduling cryo lab use with one individual using the lab per day, and announcing who will be in lab the day before.
    • If needed, this analysis suggests we could schedule two people to use the lab during the same day, but not at the same time. (of course, we are not yet cleared for shared use of the lab)
  • Use the air scrubber near the workstations, or wherever in the lab has stagnant air and prolonged use
  • Increase air circulation in the lab
    • I can order fans this week.
    • Can we open an additional port for air flow out (and possibly in) to the lab? What's behind the old cryo access in the NW corner, could we put a vent for out flow?
    • Change the HEPA filter
  • Before returning the anemometer, perhaps should measure air flow in the interior of the lab.
  • Check out the new cryo lab group and repo(s) on the github!

Cryo lab air flow

Liz dropped off an air scrubber (Medify Airx MA-40) and an anemometer (Digi-sense 20250-15) today. I'm using the instructions on the DCC to measure the air flow and assess the occupancy limits of the cryo lab. I calculate an acceptable amount of time between lab uses, and for two people to occupy the cryo lab simultaneously. For these calculations, I use a conservative threshold of P<1% for the acceptable probability that a second person becomes infected, given one infected lab occupant who sheds virus at 10 nL/min. I measured the dimensions of the lab at about 30x20x10' (l, w, h), for a 6000 ft^3 volume. I expect that's a high estimate, as it doesn't account for things like the awkward geometry of the staircase, volume of lab equipment, or stagnant air inside cabinets.

I measured the air flow at 20 points in a grid covering the surface of the HEPA filter. I also confirmed that there is no measurable air flow into the lab other than at the HEPA filter (eg, at the hole in the ceiling in the NW corner of the lab, or where the RF timing sígnal is sent through wall to Crackle lab). The average air flow through the HEPA filter is 200 CFM (cubic ft / min).
The flow is highest near the edges of the filter--does this indicate it is nearing the end of its life (it's been at least a year since replacement)? The flow is mostly above 200 CFM, but varies from 100-300 CFM across position and time.
 

Result

The first row is the minimum number of minutes to wait between when a (presumably infected) person exits the lab, and when a (presumably uninfected) person can enter the lab to maintain a <1% probability of new infection through aerosolized spread. The second row is the maximum number of person-hours the second and further (presumably uninfected) users can collectively spend in the lab with one presumably infected user, while still keeping a <1% probability of observing any areosolized spread. I don't have Spirou online currently, but I'll eventually put the calculation notebook in a shared folder accessible to the new cryo lab environment repo. Update: it's here.
 
Configuration cloth mask, no scrubber cloth mask, with scrubber N95+, no scrubber N95+,with Scrubber 
Minimum time between occupants (min) 17 7 0 0
Maximum person-hours of shared occupancy (person-hour) 0.004 0.008 0.08 0.15

These results are consistent with what I see for similar types of rooms in the LIGO spreadsheet.

Under an air flow model assuming perfect distribution of air from HVAC and HEPA scrubber throughout the lab, and very conservative requirements for probability of spreading COVID assuming a single lab user is infected, 

  1. Within 6 minutes of an infected user exiting lab, an uninfected person can enter and expect a tolerably low concentration of aerosolized virus.
  2. Within 10 minutes of one infected and one uninfected person working in the lab 6'+ apart with N95 or better masks, the viral load to an uninfected at-risk person may be intolerably high

Furthermore, the air in the lab Is not perfectly distributed. The air intake is several feet from the outflow, both near the door. The primary heat sources are the electronics racks along the N and NW sides of the room. The gradient from 71.1 F at the thermostat to 78 F at the cryocavs rack is uncomfortable to work in and bad for the electronics and optics. It's also too stagnant for the viral load conditions assumed in Evans P2000189 to apply. 

We are not currently cleared for 2+ people to use the lab. This analysis suggests to me that before doing so, we should improve the air flow conditions in the lab. And, even under optimal conditions we may minimize the total time with multiple people in the lab simultaneously.

Rio laser

I disconnected the oscillator, PDH boxes, and laser drivers from the power strip, and powered the strip with the Tenma supply. I had the ground and negative pins switched at first, careful of this -- ground is black, negative is green on this cable.

I plugged in the electronics, but only turned on DC power to the laser driver. I just left them on long enough to confirm a bright green (on viewcard) beam spot for both lasers -- we're in business!

Turned off the lasers, disinfected the common surfaces and objects.

PS, the anti-fog wipes work wonders. My goggles went from fogging in seconds to no fog at all.

  2544   Tue Jul 28 13:33:06 2020 aaronUpdateLab WorkRio laser

Notice of lab entry

[anticipated]

Date of entry: 28 July 2020
time of entry:  15:00 
time of exit:    19:30
 
Purpose: Turn on the Rio laser, find what we need for beam profiling, do some cleaning (mop, putting away stuff from the workbenches, etc)
  2543   Mon Jul 27 12:11:49 2020 aaronUpdateLab WorkSimplifying W optic bench

PSOMA hardware inventory

I located the materials for stage 1 PSOMA on the West optics table. I recorded what we have in the hardware inventory, and what we don't have is flagged for purchase. I start by cleaning up the electronics rack, removing anything I think is not in use.

Electronics rack

  • Both laser current drivers are D1500207
  • The E laser has a TED 200 C temp controller. The W laser is not plugged in to TEC, but there is a spare ITC 502 combination controller on the rack, and cryo cavs has a additional ITC 510.
  • There are three universal PDH servo boxes on the W rack: 1x D0901351 and 2x D1700219-v1. All of the boxes need lids and proper rack mount hardware.
    • Innards
      • Inside the *1351, the board matches the front panel DCC number (serial number 1437)
      • Both the *0219 have matching boards: D1700192-v2, D1700195-v2, D1799182-v2, D1700131-v2
    •  
  • Power supplies
    • The +- 15V power strip is currently unplugged. I don't see a low noise DC supply on this rack, but there is a Tenma 72-6615 on the ground. The closed-ring connector of the power strip is incompatible with this supply, so likely it was powered off of something else.
    • There's a NewFocus 1901 +- 15V supply on the W table itself. I think this should be located not on the table? The outputs are 0.1 A max to banana, plus 2x 0.3A max to a coax cable. The NewFocus was only supplying the PDs on the table. 
    • Cryo cavs has a dedicated rack for its power supplies (2x Sorensen DCS33-33E), separated from the laser control and PDH electronics. Is this an important choice, and should we also move our power supplies?
    • On the bottom of the rack, there's a Mech-Tronics NIM power supply. Seems out of use, only a PD power supply and readout board is plugged in, but this has no corresponding PD. 
    • I found on the elog that Chris moved the two Sorensens from the cryo cantilevers (W) rack to the cryoaux (central, rolling) rack after the supplies Johannes bought in 2017 died (fans not running). Do these usually die right after their 2 year warranty expires, or were we especially hard on them?
      • sidenote, great to see a potential solution to cryo lab timing woes, and that the overheated Sorensens were causing cominaux hard drive failures. Thanks Chris!
      • I've added 2x Sorensen supplies to our purchase list for PSOMA
    • For now, I moved all existing power supplies onto the electronics rack, and power the strip with the Tenma. In the future, we can discuss a remote location for power supplies, and will buy 2x additional Sorensens.
  • Misc
    • a bunch of old and open circuit boards, in cases compatible with the NIM supply. One power amp II; one PDH2 servo board (D1100996); 2 channel low noise amplifier (contains D1101396-v2 and another hand made board); a general filter board claiming to be from the gyro experiment. 
    • There were a few loose minicircuits components connected to the OCXO preamp (D1500064). These were providing 33.59 and 32.7 MHz modulation to Zach's experiments. I removed these from the preamp so we can give them a more permanent box later.

W optics bench

  • vacuum pump
    • I don't see the pumping station that used to be by the W optics bench. I don't remember using it elsewhere, and don't find any elog mentioning it being borrowed or moved. Has anyone seen this pump? It's one of the HiCube all-in-ones; it should have LIGO property tag C21832.
    • Similarly, what happened to the IR labs cryostat that housed the cantilever Q measurement (property tag)? I see maybe the glint of a cantilever in the short gryo vacuum can, and I have the cantilever Q HV driver and feedthrough, but no cryostat.
  • Photodiodes
    • PDA50B (Ge, 800-1800 nm)
    • 2x PDA255
    • Found 2 loose (unmounted) photodiodes wrapped in foil near the "optical contact razorblade test" setup (SE corner of W table).
    • NewFocus 1811, 1611
    • 1 home brewed QPD and breakout board, (PN on the circuit board is 10-00146 rev 1)
  • Put away or onto the workbench any loose optics, unused optomechanics, and most cables. Anything that wasn't a stock part I placed in a clear plastic box labeled 'Cryo Cantilevers' in the projects cabinet (2-part wooden cabinet on S wall). Stock parts went into the respective optics cabinets.

Misc

  • The lab temperature is still way too high. The temperature at the particle counter is reading 78, the thermostat still reads 71 F. The filter is not leaking substantially.j
  • Got the speaker working -- there was an aux adapter in the headphone port, so system was confused
  • Raymond requested a CF16 blank. I couldn't find one, but gave him a CF40-to-CF25 adapter that had a CF25 blank epoxied onto it, along with a CF16 to CF40 adapter and some gaskets. Here was the disinfection procedure for sending these items to QIL:
    1. Found a large plastic bag and disinfected inside and out with Phenomenal (TM) spray
    2. Disinfected the unpackaged UHV components with 70% isopropyl alcohol and a Kim wipe, wrapped them in foil, and placed in the disinfected bag
    3. Disinfect the outside of the bag with Phenomenal spray and leave outside the door to cryo for same-day pickup.
  • noted that pressure in cryo Q vacuum is 24 utorr after about a day valved off
  • Emailed Jordan and Liz about anemometer, to measure air flow in cryo. If we can compute the time for air to circulate through the cryo lab, we can reduce the length of time between lab entries (currently Shruti and I are leaving 1 day between our lab uses). 
  • Starting piling up seemingly unused components from W table onto the S workbench. These need a new home.
  • Found some historical documents, sticky notes from Zach. Took photos of these just in case (attached, later) and placed along with the miscellany on the workbench.
  • Checked with other W Bridge lab users before entering EE shop at 3:30pm. Disinfected surfaces I touched before and after use. I retrieved 3x 9V rechargeable batteries for the cryo lab multimeters.
  • Found a cardboard boxed fiber polarization controller (FPC032), which I placed in the NW optics cabinet sans box. Is this safe for storage (attachment 3)? What about things like Kim wipes, are those an acceptable amount of cardboard, or should they be eliminated?
  • Moved any cantilevers, etc into the silicon storage cabinet (NW corner cabinet)
ELOG V3.1.3-